Hi I have a question that I need help with, here is the actual question:
"The wheels on my skateboard have a diameter of 2.25 inches. If I am traveling at a rate of 15 mph, what is the angular speed of the wheels in revolutions per second?"
Now the answer I got is 37.35 revolutions/second. I don't think that is right. I was thrown off because it ask for angular speed and the formula is w=V/r, but to get revolutions wouldn't I need to divide 264in/sec by 2.25 times 2pi or just by pi?
Thanks

Answer by stanbon(48558)

"The wheels on my skateboard have a diameter of 2.25 inches. If I am traveling at a rate of 15 mph, what is the angular speed of the wheels in revolutions per second?"
Now the answer I got is 37.35 revolutions/second. I don't think that is right. I was thrown off because it ask for angular speed and the formula is w=V/r, but to get revolutions wouldn't I need to divide 264in/sec by 2.25 times 2pi or just by pi?
============
circumference of wheel = 2.25pi inches
-------------------
Distance traveled in 1 hr. = 15*5280*12 inches
-----------------------
# of revolutions in 1 hr. = (15*5280*12)/2.25pi = 134454.0959 revs/hr.
-----------------------
# of revolutions in 1 sec = (revs/hr)/(3600 sec/hr) = 37.348 rev/sec
------------------------
angular velocity = (revs/sec)*)(2pi radians) = 74.697pi or approx 234.67 radians/sec
============
Cheers,
Stan H.